To combat the current difficult situation, the manager plans to curb the variable expenses and bring them to ₹500 per piece. Compute the new break-even point and contribution margin. Analyze and explain the movement in contribution margin to the manager


1
Expert's answer
2021-11-28T17:54:27-0500

Break-even point:


"=\frac{3,000,000}{(2,500-1,000)}=2,000" units.


Contribution margin:


"=\frac{(2,500-1,000)}{2,500}=0.6"